OpenStudy (anonymous):

I've made a table with OS LaTeX and I copied and pasted it into a document on the sharelatex.com webiste. The table doesn't show up correctly on that website :(

OpenStudy (anonymous):

\[\quad \begin{array}{|c|c|c|c|} \hline \frac{\textrm{class limits}}{\textrm{lower-upper}} & \frac{\textrm{class boundary}}{\textrm{lower-upper}} & \frac{\textrm{class}}{\textrm{midpoint}} &\text{frequency}&\frac{\text{cumulative}}{\text{frequency}}&\frac{f}{n}\\ \hline 0-5 & 0.0-5.5 &13 &13& 2.75&0.218\\ \hline 6-11& 5.5-11.5& 15&28& 8.5 &0.27\\ \hline 12-17&11.5-17.5& 11&39& 14.5 &0.2\\ \hline 18-23&17.5-23.5& 3 &42& 20.5 &0.054 \\ \hline 24-29&23.5-29.5& 6 &48& 26.5 &0.109\\ \hline 30-35&29.5-35.5& 4 &52& 32.5 &0.072\\ \hline 36-41&35.5-41.5& 2 &54& 38.5 &0.036\\ \hline 42-47&41.5-47.5& 1 &55& 44.5 &0.018\\ \hline \end{array}\]

OpenStudy (anonymous):

@experimentX would you mind copying and pasting this into a sharelatex.com document and see if you get the same issue?

OpenStudy (experimentx):

do you want it in PDF?

OpenStudy (experimentx):

Woops sorry ... jpg?

OpenStudy (anonymous):

I'm not sure...I'm soo programming illiterate. I wrote my homework on sharelatex.com and this graph is just a small part of my hw assignment. I want to turn it all in together eventually.

OpenStudy (experimentx):

are you trying to change this into graph?

OpenStudy (anonymous):

Sorry I see...I worded my question wrong...hahaha I'm trying to copy and past this table into a sharelatex document. It doesn't come up as a table on sharelatex. So it's corerct on OS website but doesn't come up as a table on sharelatex.

OpenStudy (anonymous):

this is what I get...

OpenStudy (anonymous):

one issue I found is that they don't recognize the \text{} command...on their website its \textrm{}

OpenStudy (experimentx):

they all show x mark on left.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

change \text{} to \textrm{} It doesn't make too much of a difference though

OpenStudy (anonymous):

and I think they should be between $$.......$$ (dollar signs)

OpenStudy (experimentx):

he could you please send format to me ... i lost it :( messig around

OpenStudy (anonymous):

what's the format? Sorry I'm still learning the programming language... $$\quad \begin{array}{|c|c|c|c|} \hline \frac{\textrm{class limits}}{\textrm{lower-upper}} & \frac{\textrm{class boundary}}{\textrm{lower-upper}} & \frac{\textrm{class}}{\textrm{midpoint}} &\textrm{frequency}&\frac{\textrm{cumulative}}{\textrm{frequency}}&\frac{f}{n}\\ \hline 0-5 & 0.0-5.5 &13 &13& 2.75&0.218\\ \hline 6-11& 5.5-11.5& 15&28& 8.5 &0.27\\ \hline 12-17&11.5-17.5& 11&39& 14.5 &0.2\\ \hline 18-23&17.5-23.5& 3 &42& 20.5 &0.054 \\ \hline 24-29&23.5-29.5& 6 &48& 26.5 &0.109\\ \hline 30-35&29.5-35.5& 4 &52& 32.5 &0.072\\ \hline 36-41&35.5-41.5& 2 &54& 38.5 &0.036\\ \hline 42-47&41.5-47.5& 1 &55& 44.5 &0.018\\ \hline \end{array}$$

OpenStudy (experimentx):

\[ \quad \begin{array}{|c|c|c|c|} \frac{\textrm{class limits}}{\textrm{lower-upper}} & \frac{\textrm{class boundary}}{\textrm{lower-upper}} & \frac{\textrm{class}}{\textrm{midpoint}} & \textrm{frequency} & \frac{\textrm{cumulative}}{\textrm{frequency}} & \frac{f}{n} \\ 0-5 & 0.0-5.5 & 13 & 13 & 2.75 & 0.218 \\ 6-11 & 5.5-11.5 & 15 & 28 & 8.5 & 0.27 \\ 12-17 & 11.5-17.5 & 11 & 39 & 14.5 & 0.2 \\ 18-23 & 17.5-23.5 & 3 & 42 & 20.5 & 0.054 \\ 24-29 & 23.5-29.5 & 6 & 48 & 26.5 & 0.109 \\ 30-35 & 29.5-35.5 & 4 & 52 & 32.5 & 0.072 \\ 36-41 & 35.5-41.5 & 2 & 54 & 38.5 & 0.036 \\ 42-47 & 41.5-47.5 & 1 & 55 & 44.5 & 0.018 \\ \end{array} \]

OpenStudy (anonymous):

but does it show up like that on sharelatex?

OpenStudy (anonymous):

what other program can I use?

OpenStudy (experimentx):

looks like this didn't work :( try making a jgp image of code

OpenStudy (anonymous):

how would you make a table?

OpenStudy (experimentx):

if you are looking for image then it's here for you

OpenStudy (anonymous):

thanks for the image!